OpenStudy (anonymous):

Shahid spent five days hiking through the Grand Canyon. On Monday, he hiked 7.23 miles. On Tuesday, Wednesday, Thursday and Friday, Shahid hiked 4.36 miles, 5.49 miles, 8.42 miles and 6.81 miles, respectively. What is the average number of miles Shahid hiked each day? Round your answer to the nearest hundredth. A. 5.02 miles B. 6.81 miles C. 6.46 miles D. 32.31 miles

OpenStudy (lgbasallote):

\[\frac{\text{total number of miles}}{\text{number of days}}\]

OpenStudy (lgbasallote):

total -> add all of them

OpenStudy (lasttccasey):

\[\frac{7.23+5.49 +8.42 +6.81+4.36}{5}=X\]

OpenStudy (lasttccasey):

Once you add all the distances, divide by the number of trips which is 5 so X = 6.462 or "C".